OpenStudy (anonymous):

How do I explain the Pythagorean theorem in terms my teacher is capable of understanding?

OpenStudy (anonymous):

a pythagorean theorm can only be used in the triangle that have right angle, two legs and hypotenuse.. the side opposite of the 90 degree or right angle is called hypotenuse in variable = c the two other legs can be denote as a and b then the formula is c^2 = a^2 + b^2

OpenStudy (anonymous):

i know that i tutor kids who are older that me in algebra

OpenStudy (anonymous):

the problem is that my teagher doesn't understand my method of proving it but its the simplest one i can think of

OpenStudy (anonymous):

ok what is the equation you are trying to prove?

OpenStudy (anonymous):

the pythagorean theorem

OpenStudy (anonymous):

ok so a^2+b^2=c^2 ? or are you or are you trying to prove an equation using the Pythagorean theorem?

OpenStudy (anonymous):

im trying to take my proof of the pythagorean theorem and make it simpler